Two infinite parallel plates, charged with uniform surface charges densities σ1, σ2 as shown. Find the electrostatic field (in N/C) at point A: σ1 = -60 nC/m^2 , σ2 = 20 nC/m^2


1
Expert's answer
2021-01-11T11:42:16-0500

The magnitude of the field between the plates is


"E=E_1+E_2,\\\space\\
E_1=\frac{\sigma_1}{\epsilon_0},\space E_2=\frac{\sigma_2}{\epsilon_0},\\\space\\
E=\frac{1}{\epsilon_0}(\sigma_2-\sigma_1)=9035\text{ N/m}."